OpenStudy (anonymous):

A restaurant manager bought 20 packages of bagels. Some packages contained 6 bagels each and the rest contained 12 bagels each. There were 168 bagels in all. How many packages of 12 bagels did the manager buy?

OpenStudy (midhun.madhu1987):

Total number of packages of bagels purchased = 20 Let the package containing 6 bagels be x. Then the package containing 12 bagels = (20 - x) Total bagels = 168 So, 6x + 12(20 - x) = 168 Solve this to find x.

OpenStudy (anonymous):

8